# The Milky Way as Seen by Classical Cepheids I: Distances Based on Mid-infrared Photometry

### v1

@@ 3.3 Comparison to mid-IR based distances
In this paper we calculated distances to classical Cepheids using the same approach as in S19, therefore we should expect our results to be comparable. S19 provide distances to 2390 Galactic Cepheids, of which 2219 stars have new distances determined in this work. There are two reasons behind this difference. One is that the classical Cepheid list used by S19 contained stars which have since been reclassified as other types of variable stars and removed from the P21 list. The second reason is that S19 use both the Spitzer and WISE data and average individual distances from four WISE and four Spitzer bands (or as many as are available), while we use only the WISE W1 photometry, therefore some of these stars are the ones that only have mid-IR Spitzer measurements, and no WISE photometry.
The distance difference against Galactic longitude is presented in Figure 13. The overall agreement is good, as expected, especially when compared with Figure 1. There is however a mean offset of $242\pm 446$ pc which is due to several differences between these studies. One, already mentioned above, is the use of observations in all available mid-IR bands by S19. In addition, they utilize the WISE photometry from the AllWISE Multiepoch Photometry Table, which provides time-resolved observations. The second difference is that we correct for the offsets between the AllWISE and unWISE catalogs (see Section 2.2 and Eq. 3.1). This results in a large mean offset in $W1$ magnitude between this study and S19: $W1_{\rm S24}-W1_{\rm S19}=0.155\pm 0.109$ mag for $W1<8.25$ mag, while only $0.005\pm 0.045$ mag for $W1>8.25$ mag. However, when we exclude stars fainter than $W1=8.25$ mag from the distance comparison, the mean offset is still $238\pm 506$ pc, which is similar to the one for the full sample.
@@
Another major difference between S19 and this study is that we use the newer version of 3D extinction map $mwdust$ which utilizes the map of Green et al. (2019) over the map of Green et al. (2015), and a $0.470$ conversion factor between $A_{K_{S}}$ and $A_{W1}$ instead of 0.591 used by S19. For the purpose of the comparison we repeated our distance estimation procedure using $A_{W1}/A_{K_{S}}=0.591$ and Green et al. (2015) extinction maps for 1408 Cepheids with $W1>8.25$ mag. The difference in both distance and $A_{W1}$ extinction is plotted in the left and right panels of Figure 14, respectively. In blue, we show the distance/extinction difference when using the 0.470 conversion factor and Green et al. (2019) maps, while in black, with the use of 0.591 factor and Green et al. (2015) maps. As anticipated, our results are more consistent with those of S19 when using $A_{W1}/A_{K_{S}}=0.591$ and Green et al. (2015) extinction maps, the mean difference in distances being only $13$ instead of $238$ pc. We can therefore include that new distances differ from S19 primarily due to the change in the estimation of the extinction.
@@ 3.4 Comparison to distances based on the Gaia Wesenheit index
Figure 15 shows the distance difference between this study and Gaia Collaboration et al. (2023), plotted against Galactic longitude, for 2910 Cepheids that have distances from both studies. As expected, the largest discrepancies are observed toward the inner Galaxy, where the extinction is highest – the mean offset is $179$ pc (with the dispersion of $1293$) for the whole sample of 2910 stars, while it is $25$ pc (with the dispersion of $544$)for the 885 Cepheids located away from the Galactic Center ($l<-100$ or $l>100$ deg), showing that both methods give similar results in regions of low extinction. The discrepancies also increase with the $G$-band magnitude and $G-W1$ color (Figure 16). The trend of systematically larger differences with respect to color shows that extinction plays a leading role in the distance determination, as the fainter and redder stars are typically the ones that are located in regions with higher extinction.
@@ 3.5 Cepheid ages
Figure 17: Ages of 3425 Cepheids in our sample estimated from the period-age relation derived by Anderson et al. (2016). The median age is 132 Myrs.
@@ 4 Discussion
With our new distances it is worthwhile to at least make some brief remarks on the resulting spatial distribution of our Cepheids. Figure 18 shows an overdensity map of the Cepheid sample in the Galactic disk, compared to different spiral arms models available in literature. The map has been obtained with the 2-dimensional wavelet transformation (WT) code MGwave by Lucchini et al. (2023), applying the à trous algorithm (Starck & Murtagh, 1994; Starck et al., 1998) to the X-Y histogram of our sample (bin size of 75 pc). In particular, we show the wavelet scale $j=4$, which highlights mostly structures with sizes between 1.2 to 2.4 kpc. We can distinguish at least two main features emerging from the (red) overdense regions, corresponding to segments of the Perseus arm (X $\sim$ -4 kpc, Y from $\sim$ -14 to 0 kpc) and the Sagittarius-Carina arm (X $\sim$ 2 kpc, Y from $\sim$ -8 to $\sim$ 1 kpc). The comparison between the WT map and various spiral arms models available in literature indicates that the geometry proposed by Levine et al. (2006) (black dashed line in Figure 18) based on HI data is the one that better traces the continuation of the Perseus arm in the third quadrant, as already remarked by previous works (Poggio et al., 2021; Gaia Collaboration et al., 2023). On the other hand, we do not expect to see a perfect agreement between the distribution of classical Cepheids and the models based on much younger tracers, as Cepheids have already had time to migrate from their birthplaces (the regions of highest star formation), especially those in the outer disk, under the assumption that the rotational velocity of stars in the disk is different than the one of the spiral pattern. A more detailed analysis of the spiral structure using this dataset will be presented in Paper II (Drimmel et al. 2024, submitted).
@@
It is also worth noting that the vertical coordinates exhibit a clear signature of the Galactic warp, with the Cepheids being systematically shifted upwards (downwards) with respect to the Galactic plane for Ygreater-$\gtrsim$0 (Yless-$\lesssim$0), approximately up to $\sim$ 1 kpc (-1 kpc) in Z, as shown in Figure 19 (left panel). The observed warp amplitude is in agreement with the prediction from warp models based on Galactic Cepheids available in the literature (e.g. Skowron et al., 2019a, b; Lemasle et al., 2022; Cabrera-Gadea et al., 2024). The right panel of Figure 19 shows again the distribution of Cepheids in the Galactic plane, but now color-coded by age. As already remarked by previous works (e.g. Skowron et al., 2019a; Anders et al., 2024), Cepheids in the outer disk are typically older than those in the inner parts due to the Galactic metallicity gradient, because less massive (i.e., shorter period, older) stars do not reach the instability strip during their evolution, when in high-metallicity environments. However, simulations by (Skowron et al., 2019a) show that even after accounting for the metallicity gradient, we should observe more long-period Cepheids in the outer disk (farther than 16 kpc from the Galactic center), and less in regions with solar metallicity, assuming a constant star formation rate. This indicates, that such assumption might not be correct, and simulations show that in order to reproduce the current locations of Cepheids, star formation should happen in several distinct episodes in the past (for details see the supplementary material in Skowron et al. 2019a).
@@
In section 3.2 we verified our uncertainties by comparing our photometric distances to parallax-based astrometric distances, which also resulted in an estimate of the mean parallax offset (zeropoint) for the subsample of 910 Cepheids used in the analysis. Lindegren et al. (2021b, hereafter L21) characterized the parallax offset $Z_{\rm off}$ in Gaia DR3 as a function of color, magnitude and position on the sky, such that $\varpi-Z_{\rm off}$ is intended to be the true parallax (i.e. $Z_{\rm off}=-\varpi_{0}$) . If we recover their proposed offsets for our subsample of 910 Cepheids with good parallaxes, we find a mean offset $Z_{\rm off}$ of $-24.3\>\mu\rm{as}$. Our estimate of $\varpi_{0}\approx 7\>\mu\rm{as}$ suggests that the L21 offsets over-corrects the Gaia DR3 parallaxes for our sample. Meanwhile, the standard deviation of these L21 offsets is $14.7\>\mu\rm{as}$, while our measure of the dispersion of the zero point, $\sigma_{0}\approx 27\>\mu\rm{as}$, is significantly larger than this. However, our parameter $\sigma_{0}$ may also be capturing the additional uncertainty due to $\varpi_{0}$ varying on small angular scales (Lindegren et al., 2021a) as well as absorbing possible under-estimation of the parallax uncertainties at bright magnitudes (Sanders, 2023). Since the parallax offset has a complicated dependency on color, apparent magnitude and position on the sky, and is calibrated using quasars – sources that are quite different from the Cepheids – it is perhaps not surprising that our results differ from L21, who in any case encourages those who use the Gaia parallaxes to derive their own offsets appropriate for their science case. Nevertheless it is useful to compare our derived offset with other studies using Cepheids.
@@ 5 Summary
We have used both AllWISE and unWISE $W1$ photometry to estimate new distances for 3425 Galactic Classical Cepheids, which we make available in the form of a catalog. Using a subset of 910 Cepheids with good parallaxes we have confirmed that our distances are accurate with relative distance uncertainties on the order of 13%.
The accuracy of our $W1$-based distances is in part due to the relatively small extinction correction needed in the mid-IR. Indeed, for 90% of our sample the extinction is estimated to be less than 0.45 mag. In addition, we correct for extinction using a 3D extinction map, rather than using the Wesenheit index, to avoid the assumption of a universal extinction curve. In the optical this can lead to large systematic errors, while in the mid-IR imprecise colors will increase the photometric uncertainties of Wesenheit magnitudes. We instead correct for extinction using a 3D extinction map in the $K_{S}$ band and adopting an extinction ratio of $A_{W1}/A_{K_{S}}=0.47$ derived from the mid-IR extinction curve from Gordon et al. (2023). While our adopted value for $A_{W1}/A_{K_{S}}$ assumes $R(V)=3.1$, this extinction ratio varies by less than 10% from this mean value for $R(V)$ values between 2.6 to 5.5. In addition, the precision of our distances is thanks to the tight correlation of the PL diagram in the $W1$ passband, contributing only about 0.08 mag to the uncertainty in the luminosities (Wang et al., 2018).
@@
Galactic Cepheids have played an important role in understanding the structure and kinematics of the young stellar disk. In this contribution we have only briefly commented on the distribution of our sample. In an accompanying paper we will explore in more detail what this sample reveals about the spiral structure in our Galaxy (Drimmel et al. 2024, submitted). We hope that additional future Galactic studies will benefit from using these distances.
@@ Acknowledgments
DMS thanks Paweł Pietrukowicz for fruitful discussions on classical Cepheid classification. RD thanks Rene Andrae and Federico Marocco for useful conversations, Morgan Fouesneau for computing extinction ratios, and Vincenzo Ripepi and Roberto Molinaro for providing the list of sources used in Molinaro et al. (2023). We acknowledge discussions with Edward Schlafly on handling unWISE uncertainties.
@@
SK acknowledges use of the INAF PLEIADI@IRA computing resources (http://www.pleiadi.inaf.it).

### v2

@@ 3.3 Comparison to mid-IR based distances
In this paper we calculated distances to classical Cepheids using the same approach as in S19, therefore we should expect our results to be comparable. S19 provide distances to 2390 Galactic Cepheids, of which 2219 stars have new distances determined in this work, meaning that 171 stars from S19 are not in our sample. There are two reasons behind this discrepancy. One is that the classical Cepheid list used by S19 contained stars which have since been reclassified as other types of variable stars and removed from the P21 list. The second reason is that S19 use both the Spitzer and WISE data and average individual distances from four WISE and four Spitzer bands (or as many as are available), while we use only the WISE $W1$ photometry, therefore some of these stars are the ones that only have mid-IR Spitzer measurements, and no WISE photometry.
The distance difference against Galactic longitude is presented in Figure 15. The overall agreement is good, as expected, especially when compared with Figure 1. There is however a mean offset of $242\pm 446$ pc which could be due to several differences between these studies. One, already mentioned above, is the use of observations in all available mid-IR bands by S19. In addition, they utilize the WISE photometry from the AllWISE Multiepoch Photometry Table, which provides time-resolved observations. The second difference is that we correct for the offsets between the AllWISE and unWISE catalogs (see Section 2.2 and Eq. 3.1). This results in a large mean offset in $W1$ magnitude between this study and S19: $W1_{\rm S24}-W1_{\rm S19}=0.155\pm 0.109$ mag for $W1<8.25$ mag, while only $0.005\pm 0.045$ mag for $W1>8.25$ mag. However, when we exclude stars brighter than $W1=8.25$ mag from the distance comparison, the mean distance offset is still $238\pm 506$ pc, which is similar to the one for the full sample.
@@
Another major difference between S19 and this study is that we use the newer version of 3D extinction map $mwdust$ which utilizes the map of Green et al. (2019) over the map of Green et al. (2015), and a $0.470$ conversion factor between $A_{K_{S}}$ and $A_{W1}$ instead of 0.591 used by S19. For the purpose of the comparison we repeated our distance estimation procedure using $A_{W1}/A_{K_{S}}=0.591$ and Green et al. (2015) extinction maps for 1408 Cepheids with $W1>8.25$ mag. The difference in both distance and $A_{W1}$ extinction is plotted in the left and right panels of Figure 16, respectively. In blue, we show the distance/extinction difference when using the 0.470 conversion factor and Green et al. (2019) maps, while in black, with the use of 0.591 factor and Green et al. (2015) maps. As anticipated, our results are more consistent with those of S19 when using $A_{W1}/A_{K_{S}}=0.591$ and Green et al. (2015) extinction maps, the mean difference in distances being only $13$ instead of $238$ pc. We can therefore conclude that new distances differ from S19 primarily due to the change in the estimation of the extinction.
@@ 3.4 Comparison to distances based on the Gaia Wesenheit index
Figure 17 shows the distance difference between this study and Gaia Collaboration et al. (2023), plotted against Galactic longitude, for 2910 Cepheids that have distances from both studies. As expected, the largest discrepancies are observed toward the inner Galaxy, where the extinction is highest – the mean offset is $179$ pc (with the dispersion of $1293$) for the whole sample of 2910 stars, while it is $25$ pc (with the dispersion of $544$) for the 885 Cepheids located away from the Galactic Center ($l<-100^{\circ}$ or $l>100^{\circ}$), showing that both methods give similar results in regions of low extinction. However, the discrepancies between the two distance determinations become larger as stars become redder (see Figure 18), showing that these discrepancies are due to the difference in how extinction is accounted for in these two methods.
In Section 3.1.3 we demonstrated that extinction errors from $mwdust$ maps combined with those from the possible variation of $R_{V}$ between 2.6 and 3.6 could results in not more that 2.9% distance uncertainties for our sample of Cepheids, with a mean value of 1.4%. On the other hand, distance uncertainties due to $R_{V}$ deviating from the assumed value of 3.1 are much larger when optical Wesenheit magnitudes are used for distance determination. For example, in Gaia Collaboration et al. (2023), Wesenheit magnitudes from the Gaia photometry were used, according to the prescription
$W_{G}=G-R_{Gaia}(G_{BP}-G_{RP}),$ (17)
where the color coefficient $R_{Gaia}=1.9$ is a consequence of the assumed extinction law (Cardelli et al., 1989), adopting a typical value of $R_{V}=3.1$. Assuming the extinction curve of Gordon et al. (2023) instead, would only change this coefficient slightly (to about 2.0), as all recent extinction curves are very similar in the optical. However, as mentioned above, $R_{V}$ values can vary between 2.6 and 3.6 near the Galactic plane. In this range, the color coefficient $R_{Gaia}$ would vary between about 1.7 and 2.2. For a Cepheid with an observed $(G_{BP}-G_{RP})$ color of 2.5 magnitudes (the mean observed color of our Cepheid sample), the Wesenheit magnitude $W_{G}$ would then vary by ±$\pm 0.6$ magnitudes, introducing a relative distance uncertainty of 30%, which is an order of magnitude larger than the one resulting from using mid-IR based distances.
@@
As an additional test, we reversed the problem and investigated how the distance differences that we observe $(d_{S24}-d_{Gaia})$ could be due to a variation in the $R_{Gaia}$ coefficient. For the purpose of this exercise we assumed that our Cepheid distances $d_{S24}$ are correct, and for each Cepheid we calculated the value of $R_{Gaia}$ that the Cepheid should have in order for its Wesenheit-based distance $d_{Gaia}$ to match the $d_{S24}$ distance. Figure 19 shows the distribution of $R_{Gaia}$ calculated for 695 individual Cepheids within the distance bin [2, 5] kpc, so that it can be easily compared to the distribution of $R_{V}$ from Zhang & Green (2025) in the same distance bin, shown in the very bottom of their Fig. 2. While we do not expect a perfect agreement, it is readily visible that $R_{Gaia}$ follows the same trend as $R_{V}$. That is, in regions where $R_{V}$ is low (e.g. $-50^{\circ}<l<50^{\circ}$), $R_{Gaia}$ is also low, and in regions where $R_{V}$ is high (e.g. $-100^{\circ}<l<-50^{\circ}$), $R_{Gaia}$ is also high. Also, the range of variation in $R_{Gaia}$ required to account for the observed differences is reasonably modest. This is yet another argument that the variation in $R_{V}$ is the main reason behind the distance differences that we observe and further supports moving to the mid-IR regime, where the extinction is not only smaller, but also less sensitive to changes in $R_{V}$.
Figure 19: The distribution of 695 Cepheids within the distance range $2-5$ kpc in Galactic coordinates, for which both $d_{S24}$ and $d_{Gaia}$ are available. The color represents the expected value of $R_{Gaia}$ estimated from the distance difference $(d_{S24}-d_{Gaia})$ assuming that $d_{S24}$ distances are correct. To be compared with Fig. 2 from Zhang & Green (2025).
@@ 3.5 Cepheid ages
Figure 20: Ages of 3424 Cepheids in our sample estimated from the period-age relation derived by Anderson et al. (2016). The median age is 132 Myrs.
@@ 4 Discussion
With our new distances it is worthwhile to at least make some brief remarks on the resulting spatial distribution of our Cepheids. Figure 21 shows an overdensity map of the Cepheid sample in the Galactic disk, compared to different spiral arm models available in the literature, including that of Reid et al. (2019) based on masers (left panel), and that of Taylor & Cordes (1993) and Levine et al. (2006) based on the distribution of HII regions and neutral hydrogen (right panel). The background overdensity map has been obtained with the 2-dimensional wavelet transformation (WT) code MGwave by Lucchini et al. (2023), applying the à trous algorithm (Starck & Murtagh, 1994; Starck et al., 1998) to the X-Y histogram of our sample (bin size of 75 pc). In particular, we show the wavelet scale $j=4$, which highlights mostly structures with sizes between 1.2 to 2.4 kpc. We can distinguish at least two main features emerging from the (red) overdense regions, corresponding to segments of the Perseus arm (X $\sim$ -4 kpc, Y from $\sim$ -14 to 0 kpc) and the Sagittarius-Carina arm (X $\sim$ 2 kpc, Y from $\sim$ -8 to $\sim$ 1 kpc).
The comparison between the WT map and various spiral arm models available in the literature indicates that the geometry proposed by Levine et al. (2006) (black dashed line in Figure 21) based on HI data is the one that better traces the continuation of the Perseus arm in the third quadrant as traced by the Cepheids, as already remarked by previous works (Poggio et al., 2021; Gaia Collaboration et al., 2023). On the other hand, we do not necessarily expect to see a perfect agreement between the distribution of classical Cepheids and the models based on much younger tracers, such as the masers, as Cepheids have already had time to migrate from their birthplaces (the regions of highest star formation), if the angular rotation rate of stars is different than the spiral pattern speed (Skowron et al., 2019a). Indeed, the coincidence of the Cepheids and the Levine arm seen in the HI would suggest that this arm is transient. A more detailed analysis of the spiral structure using this dataset will be presented in Paper II (Drimmel et al., 2024, , in press).
It is also worth noting that the vertical coordinates exhibit a clear signature of the Galactic warp, with the Cepheids being systematically shifted upwards (downwards) with respect to the Galactic plane for Ygreater-$\gtrsim$0 (Yless-$\lesssim$0), approximately up to $\sim$ 1 kpc (-1 kpc) in Z, as shown in Figure 22 (left panel). The observed warp amplitude is in agreement with the prediction from warp models based on Galactic Cepheids available in the literature (e.g. Skowron et al., 2019a, b; Lemasle et al., 2022; Cabrera-Gadea et al., 2024; Poggio et al., 2024).
@@
The right panel of Figure 22 shows again the distribution of Cepheids in the Galactic plane, but now color-coded by age. As already remarked by previous works (e.g. Skowron et al., 2019a; Anders et al., 2024), Cepheids in the outer disk are typically older than those in the inner parts due to the Galactic metallicity gradient, because less massive (i.e., shorter period, older) stars do not reach the instability strip during their evolution, when in high-metallicity environments. However, simulations by Skowron et al. (2019a) show that even after accounting for the metallicity gradient, we should observe more long-period Cepheids in the outer disk (farther than 16 kpc from the Galactic center), and less in regions with solar metallicity, assuming a constant star formation rate. This indicates, that such assumption might not be correct, and simulations show that in order to reproduce the current locations of Cepheids, star formation should happen in several distinct episodes in the past (for details see the supplementary material in Skowron et al. 2019a).
@@
In section 3.2 we verified our uncertainties by comparing our photometric distances to parallax-based astrometric distances, which also resulted in an estimate of the mean parallax offset (zeropoint) for the subsample of 910 Cepheids used in the analysis. L21 characterized the parallax offset $Z_{\rm off}$ in Gaia DR3 as a function of color, magnitude and position on the sky, such that $\varpi-Z_{\rm off}$ is intended to be the true parallax (i.e. $Z_{\rm off}=-\varpi_{0}$ as defined in Eq. 11). If we recover their proposed offsets for our subsample of 910 Cepheids with good parallaxes, we find a mean offset $Z_{\rm off}$ of $-24.3\>\mu\rm{as}$. Our estimate of $\varpi_{0}\approx 7\>\mu\rm{as}$ suggests that the L21 offsets over-corrects the Gaia DR3 parallaxes for our sample. Meanwhile, the standard deviation of these L21 offsets is $14.7\>\mu\rm{as}$, while our measure of the dispersion of the zero point, $\sigma_{0}\approx 27\>\mu\rm{as}$, is significantly larger than this. However, our parameter $\sigma_{0}$ may also be capturing the additional uncertainty due to $\varpi_{0}$ varying on small angular scales (Lindegren et al., 2021a) as well as absorbing possible under-estimation of the parallax uncertainties at bright magnitudes (Sanders, 2023). Since the parallax offset has a complicated dependency on color, apparent magnitude and position on the sky, and is calibrated using quasars – sources that are quite different from the Cepheids – it is perhaps not surprising that our results differ from L21, who in any case encourages those who use the Gaia parallaxes to derive their own offsets appropriate for their science case. Nevertheless it is useful to compare our derived offset with other studies using Cepheids.
@@ 5 Summary
We have used both AllWISE and unWISE $W1$ photometry to estimate new distances for 3424 Galactic Classical Cepheids, which we make available in the form of a catalog. Using a subset of 910 Cepheids with good parallaxes we have confirmed that our distances are accurate with relative distance uncertainties on the order of 6%.
The accuracy of our $W1$-based distances is in part due to the relatively small extinction correction needed in the mid-IR. Indeed, for 90% of our sample the $A_{W1}$ extinction is estimated to be less than 0.45 mag, with a mean value of 0.22 mag. In addition, we correct for extinction in the mid-IR using a 3D extinction map, rather than using a Wesenheit index, to avoid systematic errors resulting from assuming a universal extinction curve. In the optical this can lead to large systematic errors in final distances, while in the mid-IR imprecise colors will increase the photometric uncertainties of Wesenheit magnitudes. We instead correct for extinction using a 3D extinction map in the $K_{S}$ band and adopting an extinction ratio of $A_{W1}/A_{K_{S}}=0.47$ derived from the mid-IR extinction curve from Gordon et al. (2023). While our adopted value for $A_{W1}/A_{K_{S}}$ assumes a ratio of total-to-selective absorption $R_{V}=3.1$, this extinction ratio varies by less than 10% from this mean value for $R_{V}$ values between 2.6 and 3.6, and would result in a relative distance uncertainty of not more than 2.1%, typically around $1\%$. In contrast, the variation of $R_{V}$ in the same range of 2.6 to 3.6 would vary the Wesenheit magnitude $W_{G}$ by ±$\pm 0.6$ mag (for the typical Cepheid in our sample), resulting in a relative distance uncertainty of 30%.
In addition, the precision of our distances is also thanks to the tight correlation of the PL diagram in the $W1$ passband, so that the scatter about the PL relation only contributes about 0.08 mag to the uncertainty in the luminosities (Wang et al., 2018).
@@
Galactic Cepheids have played an important role in understanding the structure and kinematics of the young stellar disk. In this contribution we have only briefly commented on the distribution of our sample. In an accompanying paper we will explore in more detail what this sample reveals about the spiral structure in our Galaxy (Drimmel et al., 2024, , in press). We hope that additional future Galactic studies will benefit from using these distances.
@@ Acknowledgments
We thank the Referee for useful comments that improved the manuscript, and to Robert Benjamin for providing a compilation of the Glimpse surveys. DMS thanks Paweł Pietrukowicz for fruitful discussions on classical Cepheid classification. RD thanks Rene Andrae and Federico Marocco for useful conversations, Morgan Fouesneau for computing extinction ratios, and Vincenzo Ripepi and Roberto Molinaro for providing the list of sources used in Molinaro et al. (2023). We acknowledge discussions with Edward Schlafly on handling unWISE uncertainties.
@@
SK acknowledges use of the INAF PLEIADI@IRA computing resources.